# Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation

### v3

@@ 5.1 Task Definition
We employed the following name replacement methods:
  • Cross-Language Replacement: Replacing names with uncommon names from other languages, swapping the gender in some cases (GS).
  • Same-Language Replacement: Replacing names with common names from the same language, including gender-swapped (GS) variants.
  • Anonymized Replacement: Replacing all character names with neutral labels (e.g., "Speaker 0").
  • Gender-Swap Only (GS Only): Swapping the gender of the original names without changing the language.
  • Control Tests: Introducing incorrect show names or names from other TV shows to evaluate the model’s susceptibility to data contamination (an example in  11).
5.2 Result
Our analysis focuses on understanding how different replacement strategies affect the model’s predictions. Gender-swapping and cross-cultural name replacements may disrupt the model’s memorized associations more effectively than simple anonymization. This helps us gauge the model’s inference capabilities when it cannot rely solely on pretraining knowledge. Table 6 summarizes the accuracy of GPT-4o across different replacement strategies with and without character descriptions. The "Name Return" task shows that GPT-4o achieves a perfect 100% accuracy, indicating strong memorization of original character names. However, the accuracy significantly drops for other replacement strategies, especially for cross-language and gender-swapped (GS) replacements, where the performance decreases to as low as 20.2% and 15.5%, respectively. Interestingly, anonymization (replacing all names with neutral labels like "Speaker 0" and "Speaker 1") shows a moderate increase in accuracy when descriptions are provided, suggesting the model may rely more heavily on contextual information in these cases. The results indicate that providing character descriptions can slightly improve the model’s performance, particularly in the cases of anonymized and incorrect show names replacements.
Replacement Strategy Without Description With Description
Name Return 100 100
No Replacement 52.1 47.9
Cross-Language 35.5 28.2
Cross-Language (GS) 20.2 21.8
Same Language 34.2 24.1
Same Language (GS) 15.5 16.5
Anonymized 16.7 26.4
GS Only 7.3 8.9
Wrong show names 41.5 43.6
Names from other shows 12.3 31.1
Table 6: The impact of different name replacement methods on GPT-4o’s verbatim memory ability was demonstrated. With reduced confidence in its verbatim memorization, GPT-4o is more likely to rely on reasoning. By providing descriptions, GPT-4o is more inclined to access gist memory and then apply reasoning.
5.3 Analysis
Providing character summarization evaluates two key aspects of GPT’s performance. Without summarization, the task primarily tests verbatim memory, where GPT tends to struggle. However, when character descriptions are provided, the task evaluates both memorization and inference abilities. We observe that as the task complexity increases, GPT relies more on reasoning and less on simple memorization.
6 Reasoning Validation Task
The character guessing task format is quite similar to verbatim memorization evaluation tasks, such as the name cloze task Chang et al. (2023b). We have demonstrated that tasks heavily reliant on verbatim memory can measure its reliance and help transfer the model’s usage to gist memory. Intuitively, most types of character understanding tasks should rely more on gist memory. This leads us to our third question: Can we encourage GPT to rely more on gist memory in different character understanding tasks?
@@ 6.1 Benchmark Tasks Description
Generally, character understanding tasks require the model to predict a character’s information based on a given segment of text. For script-based character understanding, the provided texts often include conversations from TV show and movie scripts, while novels typically include a mix of narrative descriptions and dialogue. The six selected benchmarks cover a range of character-centric tasks, designed to test different aspects of character understanding(Table 1):
Character Guessing This task involves identifying the speaker for each utterance in a script. Each utterance within a scene is segmented and input into the model, with the speaker’s name preceding each utterance masked and replaced by a unique special token. Consistent special tokens are used to represent the same speaker within a scene. The objective of the character guessing task is to accurately predict the identity of the speaker associated with each special token.
Coreference Resolution Given a conversation in scripts that contains multiple utterances and n character mention entity c1, c2, …, cn within it, the objective of the coreference resolution task is to assemble all mention entities that refer to the same character in a cluster.
Personality Understanding Given a text snippet describing a character from classic literature along with preceding context, the objective of the personality understanding task is to predict the most likely personality trait depicted in the snippet. The task is formulated as a multi-choice problem with distractor traits, testing the model’s ability to infer character personalities from descriptive text.
Role Detection Given dialogues and scene descriptions from scripts, the objective is to determine whether each token refers to the perpetrator. This task is framed as a multi-label classification problem, where each label corresponds to a character in the script, with a focus on detecting roles within a crime narrative context.
Open Domain Question Answering The input is a dialogue excerpt from the script and a manually crafted question, the objective of the open-domain question-answering task is to extract the specific span in the dialogue that answers the question. This can be viewed as an answer span selection task, where the model is expected to identify spans within the dialogue that contain the answer to the question.
Summarization Given dialogue utterances and scene descriptions from transcripts, the objective of the summarization task is to generate a coherent summary that captures key plot points while excluding irrelevant details.
6.2 Perturbed Method
We implemented the name-changing strategy of swapping genders across cultures, which, as demonstrated in Section s 5, significantly drops performance.
6.3 Result
Table 7 presents the impact of replacing character names on a variety of downstream tasks using the GPT-4o model. The experiments assess performance across six tasks:
Memory-Driven Tasks:
These tasks rely heavily on verbatim memory and are significantly affected by name replacement.
Intuively, TV Show Guessing and Screenplay Coreference should rely more on verbatim memory and they drop of 14.6% and 8.7% meets our expectation.
Reasoning-Driven Tasks:
These tasks primarily rely on gist memory and are more likely to engage in reasoning without external hints(CoT or extra description).
The slight drops in FriendsQA and Personet personality understanding tasks, with a decrease of 0.7% and 2%, align with our expectations. The 17.2% drop in role extraction is unexpectedly high, resembling the trend seen in coreference resolution tasks. The drop in ScreenSum is moderate, falling within expected ranges.
Figure 2: All tasks exhibited performance degradation. The decline in QA and personality understanding tasks was minimal, indicating that GPT effectively leveraged its reasoning abilities for these tasks. The most significant drop was observed in role detection, exceeding our expectations, which suggests that verbatim memory plays a substantial role in supporting this type of task.
Task Name Origin Script Name Replaced
TVShow Guessing 44.8 30.2
Screenplay Coreference 58.3 49.6
FriendsQA 29.3 28.6
ScreenSum (ROUGE_L) 19.1 13.8
CSI Role Detection 50.1 32.9
PERSONET Personality 40.0 38.0
Table 7: All tasks showed a decline in performance, with the smallest drop in QA, which exceeded our expectations. We believe this is because the questions asked were relatively simple and basic. There was almost no change in personality understanding, indicating that the model knows it should use gist memory for such tasks. The largest decline was observed in role extraction, which also exceeded our expectations, suggesting that the model heavily relies on verbatim memory when handling this type of task.
6.4 Analysis
Our experiment, through name replacement, increases the difficulty of the model relying on verbatim memory. As a result, the model is more likely to focus on gist memory, thereby enabling reasoning. By comparing the performance before and after name replacement, we can observe whether the model relies more on verbatim or gist memory without external hints. Specifically, if the performance difference is small, it suggests that the model is more likely to use reasoning for answering such tasks. Conversely, if the difference is large, it indicates that the model tends to rely more on verbatim memory. All tasks showed a decline, and we observed that the model always tends to prioritize verbatim memory. This ’lazy’ behavior suggests that to unlock the model’s true reasoning ability, we need to design methods that encourage the model to focus more on gist memory, such as reducing the similarity between the text and pretraining data.
@@ 7 Conclusion
In this paper, we evaluate the extent of GPT-4’s verbatim memory for fictional works and its impact on downstream tasks. We also propose methods to mitigate this impact and encourage the model to use reasoning. We draw several key conclusions: (1) The model always tends to first rely on verbatim memory when performing character understanding tasks, and it consistently benefits from this approach. (2) When the predicted text differs significantly from the pretraining data, the model is more likely to use reasoning to attempt to compensate.
@@ B.1 Source Prediction
Appendix C Self-generated Character Description
For the self-generated character description, we use following prompts to ask ChatGPT to generate a brief character summarization for each character. The following block-quoted text shows the output we got; importantly, it demonstrates the level of deep knowledge about each of these characters that is baked into the model itself.
Input: Give me a detailed character description for all main characters in the TV show Friends.
Output: Rachel Green (Jennifer Aniston):
Background: Initially introduced as a runaway bride, Rachel starts as a waitress at Central Perk and evolves into a successful fashion executive.
Character Traits: Fashion-conscious, independent, initially somewhat spoiled, but grows into a more self-sufficient and career-focused woman.
Romantic Arc: Has an iconic on-again, off-again relationship with Ross throughout the series.
Ross Geller (David Schwimmer):
Profession: Paleontologist working at a museum.
Character Traits: Intelligent, nerdy, and somewhat socially awkward. Ross has a complicated love life and is often portrayed as the "smart" but emotionally vulnerable member of the group.
Romantic Arc: His tumultuous relationships, particularly with Rachel, are central to the series.
Monica Geller (Courteney Cox):
Profession: Chef and later a restaurant owner.
Character Traits: Competitive, organized, and slightly obsessive about cleanliness. Monica is Ross’s sister and becomes the motherly figure of the group.
Romantic Arc: Her relationship with Chandler Bing becomes a major storyline.
Chandler Bing (Matthew Perry):
Profession: Holds an office job with a mysterious title related to statistical analysis.
Character Traits: Known for his sarcastic humor, commitment issues, and use of humor as a defense mechanism. Evolves into a more mature and emotionally secure character.
Romantic Arc: Develops a significant romantic relationship with Monica.
Joey Tribbiani (Matt LeBlanc):
Profession: Aspiring actor throughout the series.
Character Traits: Easygoing, simple-minded, and often portrayed as the least academically inclined. Known for his catchphrase "How you doin’?" and his love for food.
Romantic Arc: Has numerous short-term relationships, and his genuine friendship with Chandler is a key dynamic.
Phoebe Buffay (Lisa Kudrow):
Profession: Eccentric musician with a unique and mysterious past.
Character Traits: Quirky, free-spirited, and spiritually inclined. Phoebe often provides offbeat humor and memorable songs.
Romantic Arc: Her romantic life includes unique relationships, and her eccentricity adds a distinctive element to the group.
These six characters form the core group of friends in the series, and their individual personalities, relationships, and growth over the ten seasons contribute to the show’s humor, heart, and enduring popularity.
@@ Appendix D Prompts for GPT4
for personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’
@@ Appendix G Benchmark Datasets
  • TVShow Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al. (2022c), this is a multi-choice classification task. The dialogue is anonymized using placeholder identifiers (e.g., P1, P2), and the model must predict the correct speaker for each line, testing its ability to identify characters based on context.

### v4

@@ 5.1 Name Replacement Strategies and Their Impact on Memorization
Previous studies Yu et al. (2022) have primarily employed Same-Cultural Name Replacement, a naive and intuitive approach that substitutes character names with culturally consistent alternatives. To systematically analyze the impact of name replacement on model performance, we introduce three distinct levels of name modifications:
  • Name Masking: Replacing all character names with placeholders (e.g., Sheldon → P0, Amy → P1).
  • Cross-Cultural Name Replacement: Substituting names with those from a different cultural background while maintaining gender consistency (e.g., SheldonJie Zhang, AmyXiaoling Wang). This tests whether models rely on specific name distributions for character inference.
  • Same-Cultural Name Replacement: Replacing names with other culturally consistent alternatives (e.g., SheldonAndrew, AmySally). This retains cultural familiarity while disrupting direct memorization.
Memorization Evaluation Task Format
We design the Source Identification Task where the model is given a dialogue excerpt without source annotations and must predict the correct origin of the text. Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where each utterance $U_{i}$ is part of a larger script from a specific source. The task requires the model to determine the correct source, $f_{\text{source}}(D)\rightarrow S$, where $S$ denotes the source work. We use the Fictional Works Memorization Benchmark Dataset described in Section 3.2.
Impact of Name Replacement on Memorization.
According to Figure 3, our results demonstrate that different name replacement strategies have a significant impact on the memorization ability of LLMs. Among them, Cross-Cultural Name Replacement consistently induces the strongest forgetting effect across all models, indicating that altering names to those from a different cultural background disrupts memorized associations more effectively than other replacement methods.
Figure 3: The heatmap presents accuracy percentages for source identification across three LLMs under different name replacement conditions. All three models demonstrate strong memorization of the script. The Cross-Culture Name Replacement method exhibits a significantly strong and the most pronounced memorization-weakening effect across all three models.
5.2 Memorization-Controlled Prompting under Name Replacement
Different name replacement strategies have shown a significant impact on the memorization ability of LLMs. Building on this, we apply the name replacement strategies introduced in Section 5.1 to evaluate the Character Guessing task using the different prompting strategies discussed in Section 4.2.
Under the Name Replacement condition, we provide brief character descriptions generated by GPT-4, aiming to enable the model to complete the Character Guessing task through character matching. The descriptions used are provided in the Appendix D. To assess the actual effectiveness of these descriptions, we conduct an ablation study.
Figure 4: The line chart shows the decline in memorization across different settings, illustrating that as memorization weakens, overall model performance also decreases. The bar chart presents the effect of different prompting strategies: under Name Replacement conditions, Verbatim prompting consistently outperforms the baseline by reinforcing memorization, whereas Gist prompting further reduces reliance on memorization, leading to the lowest accuracy.
Results and Key Findings
We present the results in Figure 4 and summarize our main findings as follows:
  1. 1.
    Strong Correlation Between Memorization and Performance. Model performance aligns closely with memorization levels, indicating a strong dependence on memorization for accurate predictions.
  2. 2.
    Minimal Impact of Character Descriptions . As an ablation study, we evaluate the effect of character descriptions under the Origin Script condition. Our results show that incorporating character descriptions does not significantly affect model performance, ruling out potential interference from inaccurate or misleading descriptions.
  3. 3.
    Verbatim Prompting Enhances Memorization. Across all three name replacement strategies, Verbatim prompting consistently improves model performance over the baseline, demonstrating that prompting can reinforce memorization and consequently boost accuracy.
  4. 4.
    Gist Prompting Reduces Reliance on Memorization. Performance under the Gist condition declines compared to the baseline, confirming that our Gist prompt effectively reduces the model’s reliance on memorization.
  5. 5.
    Name Replacement + Gist Prompting as an Effective Evaluation Strategy. We conclude that combining Cross-Culture Name Replacement with Gist prompting provides a reasonable and effective method for evaluating a model’s true capabilities while minimizing reliance on memorization.
6 Evaluating Character Understanding Tasks Under Memorization Constraints
In this section, we evaluate different types of character understanding tasks using the Gist Prompting with Cross Culture Name Replacement approach. We first introduce the design principles for Gist Prompts across various task types, followed by an analysis of the results obtained from three different LLMs.
@@ 6.1 Benchmark Tasks Description
We define the five character understanding tasks being benchmarked and provide an illustrative example.
Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where $U_{i}$ is an utterance in the conversation. Each utterance $U_{i}$ is associated with a speaker label $S_{i}$, and additional annotations depending on the task.
  • Coreference Resolution: Cluster character mentions $C=\{c_{1},c_{2},...,c_{n}\}$ into groups referring to the same entity, $f_{\text{coref}}(C)\rightarrow\{\text{clusters of }c_{i}\}$.
  • Personality Understanding: Classify a character’s personality trait $P$ given dialogue context $D$, $f_{\text{personality}}(D,S_{i})\rightarrow P$.
  • Role Detection: Determine whether a character plays a key role (e.g., perpetrator in a crime narrative), $f_{\text{role}}(S_{i})\rightarrow\{0,1\}$.
  • Open-Domain Question Answering: Extract the span $A$ in $D$ that answers a given question $Q$, $f_{\text{QA}}(D,Q)\rightarrow A$.
  • Summarization: Generate a summary $S$ that captures key events and relationships, $f_{\text{summ}}(D)\rightarrow S$.
Example Dialogue
To illustrate these tasks, consider the following conversation:
P0: "You’re still going to have to convince a jury that I killed two strangers for no reason."
Grissom doesn’t look worried. He takes his gloves off and puts them on the table.
P1: "You ever been to the theater, Peter? There’s a play called Six Degrees of Separation."
The correct speaker labels are P0 → Peter Berglund, P1 → Grissom. For all tasks except Character Guess, the placeholders P0 and P1 are replaced with their original names, Peter Berglund and Grissom. Table 2 illustrates how this dialogue segment is utilized across different tasks.
Task Example Question and Expected Output
Coreference Resolution "Which entity does ’his’ refer to?" → "Grissom"
Personality Understanding "What trait best describes P1?" → Analytical
Role Detection "Identify all entities referring to the criminal." → Peter Berglund
Question Answering "What play does Grissom mention?" → "Six Degrees of Separation"
Summarization "Generate a summary." → "Peter discusses convincing a jury, while Grissom references a play."
Table 2: Examples of different character understanding tasks applied to the same dialogue.
6.2 Gist Prompts Across Character Understanding Tasks
General Guidelines for Gist Prompting
  • Encourage Structural Reasoning: Guide models to infer character identity and relationships from discourse patterns rather than relying on explicit name associations.
  • Eliminate Memorization Bias: Ensure prompts emphasize evidence-based inference rather than knowledge retrieval.
  • Preserve Narrative Coherence: In summarization and question answering, focus on maintaining logical flow rather than phrase-matching.
Detailed gist prompts are included in Appendix H.
{NiceTabular}
l|ccc|ccc|ccc Tasks GPT-4o LLaMA3.3-70B DeepSeek V3
Origin NR NR+GIST Origin NR NR+GIST Origin NR NR+GIST
Tvshow Guess 78.2 32.0 $\downarrow$46.2 31.5 $\downarrow$0.5 61.4 31.9 $\downarrow$29.5 27.8 $\downarrow$4.1 60.8 29.7 $\downarrow$31.1 26.7 $\downarrow$3.0
Coreference 58.3 49.6 $\downarrow$8.7 47.5 $\downarrow$2.1 56.7 48.3 $\downarrow$8.4 48.1 $\downarrow$0.2 47.4 34.6 $\downarrow$12.8 32.3 $\downarrow$2.3
FriendsQA 44.3 38.6 $\downarrow$5.7 38.6 $\uparrow$0.0 45.7 42.9 $\downarrow$2.8 44.3 $\uparrow$1.4 42.8 39.4 $\downarrow$3.4 34.3 $\downarrow$5.1
ScreenSum 37.5 28.7 $\downarrow$8.8 23.1 $\downarrow$5.6 40.1 25.9 $\downarrow$14.2 21.3 $\downarrow$4.6 35.6 19.3 $\downarrow$16.3 17.1 $\downarrow$2.2
CSI Role Extract 50.1 32.9 $\downarrow$17.2 31.7 $\downarrow$1.2 48.3 34.7 $\downarrow$13.6 32.7 $\downarrow$2.0 45.1 29.8 $\downarrow$15.3 23.7 $\downarrow$6.1
PERSONET 56.0 52.0 $\downarrow$4.0 54.0 $\uparrow$2.0 62.0 46.0 $\downarrow$16.0 44.0 $\downarrow$2.0 44.0 40.0 $\downarrow$4.0 38.0 $\downarrow$2.0
Table 3: Ablation study examining the impact of name replacement and the GIST method on model performance across six NLP tasks.The subscripted values in the NR column indicate the performance drop relative to the original script, while the subscripted values in the NR+GIST column show the performance change compared to the Name Replace condition.
6.3 Results and Key Findings
Our main results and findings are summarized as follows:
  1. 1.
    Significant Performance Drop with Name Replacement: Across all tasks, we observe a substantial decline in performance when applying the name replacement (NR) method in Table 6.2. This result indicates that reducing explicit memorization has a strong impact on model performance, confirming our hypothesis that existing character understanding tasks rely heavily on memorization. Moreover, it highlights the potential inflation of benchmark scores under standard evaluation settings.
  2. 2.
    Further Degradation with Gist Prompting: The additional performance drop induced by Gist prompting demonstrates its effectiveness across various character understanding tasks. This suggests that Gist prompting can further encourage models to rely less on memorization, reinforcing our findings on the role of memorization in these tasks.
  3. 3.
    Memorization Can Be Detrimental to Character Understanding: Interestingly, we observe performance gains in certain cases under the combined NR+Gist setting. Our error case analysis reveals that these improvements stem from resolving knowledge conflicts—instances where the original memorization-based responses misled the model, whereas reducing memorization mitigated such errors, leading to better predictions.
  4. 4.
    Consistent Trends Across Models: As shown in Figure 5, despite variations in absolute performance, different models exhibit similar trends in performance degradation across all tasks. This consistency suggests that various task designs interact similarly with memorization, reinforcing the broader implications of our findings.
Figure 5: Performance degradation across LLMs due to the effects of GIST+NR perturbations.
@@ 7 Conclusion
In this work, we introduce the concept of gist memory and design prompts for six types of character understanding tasks, enabling models to reduce reliance on memorization and enhance reasoning-based character interpretation. Our approach provides a systematic framework to evaluate the extent to which different character understanding tasks depend on memorization, offering a more accurate assessment of model performance by disentangling genuine reasoning from memorization-driven responses. These findings highlight the limitations of current benchmarks and provide insights for designing more robust character understanding evaluations.
@@ B.1 Source Prediction
Appendix C Extended Dataset Analysis
Task Name Task Type Source
Character Mining Character Identification TV Shows
LISCU Character Identification Book Reviews
MovieQA Visual Question Answering Movies
Liter Coreference Character Coreference Novels
HVV Detect Role Detection Movies, News
SYMON Visual Question Answering Movies
CEB Character Representation Novels
Love and Violence Summarization TV Shows
@@ Appendix D Prompts for GPT4
for personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’ Character Description Here is a character description:Monica: Responsible, caring, and organized. Assertive and confident in her actions.Chandler: Witty and self-deprecating with an approachable sense of humor. Exhibits insecurity and anxiety in his dialogues, making references to uncomfortable situations and questioning his own actions.Rachel: Spontaneous and open to change, she takes risks and is adaptable. She is also reliant on her relationships with others, signifying her dependency and need for support.’ ’Joey: Energetic, extroverted, and casual. Lacks the sensitivity of others feelings at times but genuinely care about friends.Phoebe: Quirky, eccentric, and a free spirit. Her train of thought tends to lean toward the unusual and bizarre. However, she is also compassionate and caring.Ross: Insecure and somewhat neurotic and vulnerable. His behavior is indicative of someone who is going through emotional turmoil.’
@@ Appendix G Benchmark Datasets
  • TVShow Character Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al. (2022c), this is a multi-choice classification task. The dialogue is anonymized using placeholder identifiers (e.g., P1, P2), and the model must predict the correct speaker for each line, testing its ability to identify characters based on context.
@@
Appendix H Gist Prompting Strategies Across Tasks
We design Gist Prompting methods to reduce reliance on memorization and promote inference-based character understanding. Below, we detail the task-specific prompt formulations.
H.1 Task-Specific Gist Prompts
Character Guessing
"Analyze the dialogue and infer the most likely speaker based on their speech patterns, personality traits, and relationships with other characters. Do not rely on direct recall of specific names but rather use reasoning from the provided context."
Coreference Resolution
"Identify which mentions in the text refer to the same character by analyzing linguistic patterns, contextual cues, and discourse coherence. Do not rely on pre-learned entity associations but rather resolve coreference based on narrative structure."
Personality Understanding
"Based on the character’s dialogue and actions, infer their personality traits. Focus on how they interact with others, their tone, and behavioral tendencies. Avoid using prior memorized descriptions of characters and derive insights purely from the given passage."
Role Detection
"Determine whether a character plays a key role in the narrative (e.g., antagonist, victim, or protagonist) by analyzing their function within the story. Identify how their actions and relationships define their role rather than relying on predefined labels."
Open-Domain QA
"Answer the question using only the given passage. Identify relevant information by reasoning over the text and analyzing interactions between characters. Do not rely on external knowledge or prior memorized facts."
Summarization
"Summarize the passage by identifying key events, character interactions, and narrative progression. Construct a coherent summary that captures the essence of the scene without directly copying text segments."